# ImpedanceGPT: VLM-driven Impedance Control of Swarm of Mini-drones for Intelligent Navigation in Dynamic Environment

### v1

@@ IV Experimental Setup
In total, seven experiments were performed to evaluate the efficacy of the proposed ImpedanceGPT framework for semantic understanding, obstacle identification, and intelligent navigation in both static and dynamic environments, ranging from sparse to dense scenarios. These experiments also included unseen dynamic scenarios, for which no pre-existing database was available, to further assess the framework’s robustness in handling novel environments. The top-down images for each experiment were captured in daylight conditions to ensure consistent lighting and reliable image analysis. Multi-colored obstacles were strategically placed within the environment to test the semantic classification capabilities of the VLM model and to check for any potential bias.
The VLM-RAG framework was executed on a high-performance PC equipped with an RTX 4090 graphics card (24GB VRAM) and an Intel Core i9-13900K processor to ensure smooth and efficient model inference. Due to memory constraints, the quantized Molmo-7B-D BnB 4-bit model [19] was used, which provided a good balance between performance and memory usage.
For the real-world implementation of the system, Crazyflie 2.1 drones were utilized, allowing practical testing in dynamic environments. The VLM-RAG system was given the following text prompt for obstacle detection and information retrieval: “Analyze the drone arena image and identify all tripod stands or humans as obstacles in the scene. Count the total number of tripod stands or humans, specifying the number of obstacles before the gate and the number of obstacles after the gate. If there are no obstacles before the gate, mention that explicitly. Calculate the relative distances between each obstacle and describe their spacing as ‘closely spaced’ or ‘widely spaced.’ The spacing should be based on the distance between their feet in the image”.
To ensure safety during real-world experiments with humans as obstacles, participants wore thick jackets and goggles for additional protection. This setup allowed for a thorough evaluation of the system’s obstacle detection and classification capabilities in dynamic real-world environments.
@@ V-A Evaluation of VLM-RAG Framework
The experimental setups in this work represent an indoor environment, where lighting plays a crucial role in the performance of the VLM-RAG system.
Figure 4: Performance of VLM-RAG system under different lighting conditions.
Fig. 4 shows that the VLM-RAG system has a success rate of 80% in both the detection and the retrieval of soft and hard objects under optimal lighting conditions. However,in inadequate lighting, it struggles to identify difficult obstacles and thus its success rate drops to 60%. The results are presented in Fig. 4 are averaged over multiple scenarios tested in our experimental setup. Although lighting conditions affect system performance, the model successfully detects obstacles and determines their positions regardless of their color or placement.
@@ V-B Results in Static Environment
Figure 5: Scenario 1 - Environment with 4 static hard obstacles and a rectangular gate.
Figure 6: Scenario 2 - Environment with 2 static soft obstacles and a rectangular gate.
@@
  • Scenario 1 - Static: This scenario consists of four hard obstacles with a rectangular gate positioned within the environment. The drones are required to navigate through the space while avoiding collisions and reaching the designated goal position. As shown in Fig. 5, the drones successfully completed the task without any collisions. It is important to note that the overlapping trajectories of the follower drones do not indicate collisions but rather represent their positions at different time steps.
  • Scenario 2 - Static: In this scenario, two individuals (soft obstacles) are positioned before a rectangular gate in a way that obstructs the swarm of drones’ path. Despite this, the drones successfully maneuver around the human obstacles and proceed without any incident as illustrated in Fig. 6.
A key observation from these experiments is the effect of obstacle type on drone behavior. In Scenario 1, where only hard obstacles are present, the trajectories overlap significantly, indicating higher movement speeds. Conversely, in Scenario 2, where soft obstacles (humans) are introduced, the drones maintain a greater separation distance, adjusting their paths based on the optimal parameters for impedance control from the ImpedanceGPT framework. Additionally, the maximum velocity recorded in Scenario 1 is 1.4 m/s, whereas in Scenario 2, the velocity is reduced to 0.7 m/s, reinforcing the system’s ability to ensure safe navigation by modulating speed according to obstacle type.
@@ V-C Results in Dynamic Environment
Experiments in dynamic environments introduced additional complexities, including moving obstacles and changing goal positions, to assess the adaptability of the drone swarm.
Figure 7: Scenario 3 - Environment with 2 hard obstacles, a gate and moving goal position.
Figure 8: Scenario 4 - Environment with 2 soft obstacles; one was moving and the other was stationary.
Two such dynamic scenarios are detailed below:
  • Scenario 3 - Dynamic: This scenario involves two follower drones guided by a leader drone, which actively tracks a moving goal position. The drones successfully avoid obstacles while continuously adjusting their paths in response to environmental changes. As illustrated in Fig. 7, the leader drone and its followers dynamically navigate the environment, maintaining safe distances while ensuring smooth and adaptive motion toward the goal.
  • Scenario 4 - Dynamic: In this scenario, two humans serve as obstacles. One individual remains stationary in the path of the swarm, while the other moves freely. The swarm of drones successfully avoids the moving person and navigates past the stationary human, maintaining a safe distance and reducing speed as necessary to ensure collision-free navigation, as seen in Fig. 8.
A comparison of these dynamic scenarios further highlights the system’s ability to adapt motion based on obstacle type. In Scenario 3, where drones navigate around a hard obstacle, they maintain a velocity of 1 m/s, enabling efficient movement through the environment. In contrast, in Scenario 4, where a human is dynamically moving, the velocity is further reduced to 0.6 m/s, demonstrating the system’s responsiveness in ensuring safe navigation by modulating speed when encountering soft, dynamic obstacles. This adaptive behavior is crucial for maintaining safety and efficiency in real-world swarm navigation.
@@ VI Conclusion and Future Work
This paper presented ImpedanceGPT, a VLM-RAG-driven framework designed for adaptive swarm navigation in dynamic environments. Through a series of experiments, the system demonstrated its capability in obstacle classification, impedance-based motion adaptation, and intelligent navigation.
The key contributions of this work are as follows:
  • The VLM-RAG framework achieved a success rate of 80% under optimal lighting.
  • In both static and dynamic environments, drones exhibited higher velocities for hard obstacles i.e., 1.4 m/s and 1.0 m/s, respectively, while their velocities drop down to 0.7 m/s and 0.6 m/s, respectively for soft obstacles, ensuring safe and adaptive navigation.
Future work will focus on enhancing generalization through transfer learning, improving real-time adaptability via multi-objective optimization, incorporating edge computing for reduced latency, and advancing human-swarm interaction through natural language processing and gesture recognition. By addressing these areas, ImpedanceGPT aims to further improve the autonomy, adaptability, and safety of drone swarms in real-world applications.

### v2

@@ IV Experimental Setup
In total, seven experiments were performed to evaluate the efficacy of the proposed ImpedanceGPT framework for the semantic understanding, obstacle identification, and intelligent navigation in both static and dynamic along with sparse and cluttered environments. These experiments also covered unseen dynamic scenarios, for which no pre-existing database was available, to further assess the framework’s robustness. The top-down images for each experiment were captured in daylight conditions to ensure consistent lighting and reliable image analysis. Multi-colored obstacles were placed within the environment to test the semantic classification capabilities of the VLM model and to check for any potential bias.
The VLM-RAG framework was executed on a high-performance PC equipped with an RTX 4090 graphics card (24GB VRAM) and an Intel Core i9-13900K processor to ensure smooth and efficient model inference. Due to memory constraints, the quantized Molmo-7B-O BnB 4-bit model was used. For the real-world implementation of the system, Crazyflie 2.1 drones were utilized, allowing practical testing in dynamic environments. The VLM-RAG system was given the following text prompt for obstacle detection and information retrieval: “Analyze the drone arena image and identify all cylindrical stands or humans as obstacles in the scene. Count the total number of cylindrical stands or humans, specifying the number of obstacles before the gate and the number of obstacles after the gate. If there are no obstacles before the gate, mention that explicitly. Calculate the relative distances between each obstacle and describe their spacing as ‘closely spaced’ or ‘widely spaced.’ The spacing should be based on the distance between their feet in the image.” The spacing between obstacles is determined based on the Euclidean distance computed from the normalized coordinates provided by Molmo. If the distance exceeds 20 grid cells, the obstacles are categorized as widely spaced; otherwise, they are considered closely spaced.
@@ V-A Evaluation of VLM-RAG Framework
The experimental setup represents an indoor environment, where lighting plays a crucial role in the performance of the VLM-RAG system.
Figure 3: Performance of VLM-RAG system under varying lighting conditions. Optimal lighting denotes bright and well-illuminated environments, while inadequate lighting corresponds to dim or poorly lit scenarios.
A total of seven experiments were conducted, with each experiment repeated three times. The VLM-RAG system achieved an 80% success rate in both the detection and retrieval of soft and hard obstacles under optimal lighting conditions. However, under inadequate lighting, the system struggled to identify obstacles of complex shape and close proximity, resulting in a reduced success rate of 60%, as shown in Fig. 3. The results are averaged across multiple scenarios in our experimental setup. While lighting conditions do affect performance, the model consistently detects obstacles and determines their positions regardless of color or placement. The computational time for each experiment ranged from 5 to 8 s.
@@ V-B Results in Static Environment
Figure 4: Scenario 1. Environment with 4 static hard obstacles and a rectangular gate.
Figure 5: Scenario 2. Environment with 2 static soft obstacles and a rectangular gate.
@@
  • Scenario 1, Static: This scenario consists of four hard obstacles with a rectangular gate positioned within the environment. The drones are required to navigate through the space while avoiding collisions and reaching the designated goal position. As shown in Fig. 4, the drones successfully completed the task without any collisions. Overlapping trajectories of the follower drones do not indicate collisions but represent their positions at different time steps.
  • Scenario 2, Static: In this scenario, two individuals (soft obstacles) are positioned before a rectangular gate in a way that they obstruct the path of swarm. Despite this, the drones successfully maneuver around the humans and proceed without any incident (see Fig. 5).
A key observation from these experiments is the effect of obstacle type on the drone behavior. In Scenario 1, where only hard obstacles are present, the trajectories overlap significantly, indicating higher movement speeds. Conversely, in Scenario 2, where soft obstacles (humans) are introduced, the drones maintain a greater separation distance, adjusting their paths based on the optimal parameters for impedance control generated by ImpedanceGPT. Additionally, the maximum velocity recorded in Scenario 1 is 1.4 m/s, whereas in Scenario 2, the velocity is reduced to 0.7 m/s, reinforcing the system’s ability to ensure safe navigation by modulating speed according to obstacle type.
@@ V-C Results in Dynamic Environment
Experiments in dynamic environments introduced additional complexities, including moving obstacles and changing goal positions, to assess the adaptability of the drone swarm. Two such dynamic scenarios are specified below:
Figure 6: Scenario 3. Environment with 2 hard obstacles, a gate and moving goal position.
Figure 7: Scenario 4. Environment with 2 soft obstacles, one moving and one stationary.
  • Scenario 3, Dynamic: This scenario involves two follower drones guided by the leader drone, which actively tracks a moving goal position. The drones successfully avoid obstacles while continuously adjusting their paths in response to the environmental changes. As illustrated in Fig. 6, the leader drone and its followers dynamically navigate the environment, maintaining safe distances while ensuring smooth and adaptive motion towards the goal.
  • Scenario 4, Dynamic: In this scenario, two humans serve as obstacles. One individual almost remains stationary in the path of the swarm, while the other moves freely. The swarm of drones successfully avoids the moving person and navigates past the stationary human, maintaining a safe distance and reducing speed as necessary to ensure collision-free navigation, as seen in Fig. 7.
A comparison of these dynamic scenarios further highlights the system’s ability to adapt motion based on the obstacle type. In Scenario 3, where drones navigate around a hard obstacle, they maintain a velocity of 1 m/s, enabling prompt movement through the environment. In contrast, in Scenario 4, where a human is dynamically moving, the velocity is further reduced to 0.6 m/s, demonstrating the system’s responsiveness in ensuring safe navigation by modulating speed when encountering soft dynamic obstacles. This adaptive behavior is crucial for maintaining safety and efficiency in real-world swarm navigation.
V-D Limitations
The operational workspace is directly influenced by the field of view (FOV), where larger fields enable broader environments and more complex mission scenarios. In our case, the effective FOV of the available camera was limited to a $2\times 2$ meter grid, which restricted the spatial scale of the experiments. Despite this constraint, we designed and executed scenarios of increasing complexity while ensuring safe and reliable UAV navigation within the available space.
@@ VI Conclusion and Future Work
We have developed ImpedanceGPT, a VLM-RAG-driven impedance control designed for adaptive swarm navigation in dynamic environments. Through a series of experiments, the system demonstrated its capability in obstacle classification, impedance-based motion adaptation, and intelligent navigation. The VLM-RAG framework achieved a success rate of 80% in optimal lighting condition. In both static and dynamic environments, drones exhibited higher velocities for hard obstacles, i.e., 1.4 m/s and 1.0 m/s, respectively, while for soft obstacles, their velocities dropped down to 0.7 m/s and 0.6 m/s, respectively, thus ensuring safe and adaptive navigation.
Future work will focus on enhancing generalization by transfer learning, improving real-time adaptability via multi-objective optimization, incorporating edge computing for reduced latency, and advancing human-swarm interaction through natural language processing and gesture recognition.
Acknowledgements
Research reported in this publication was financially supported by the RSF grant No. 24-41-02039.